complete the ordered pairs so that each is a solution for the given equation.
3x+4y=12 (0, ), (,3/4), (,0), (8/3,)

Answer by checkley71(8403)

3X+4Y=12 X=O Y=3 Y=3/4 X=3 Y=0 X=4 X=8/3 Y=1